40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 3 22:37:44 2012, Jenne, Update, General, OpLev 90-day trend OpLevTrends_90days_Ending3May2012
    Reply  Fri May 4 17:44:42 2012, Jenne, Update, General, OpLev 90-day trend 
Message ID: 6602     Entry time: Fri May 4 17:44:42 2012     In reply to: 6601
Author: Jenne 
Type: Update 
Category: General 
Subject: OpLev 90-day trend 

Quote:

After fixing the PRM tonight, all of the oplevs look okay. 

.....except ITMX, which looks like it's power dropped significantly after the CDS upgrade.  To be investigated tomorrow.

 I had a look-see at ITMX's oplev.  I can't see any clipping, so maybe the power is just low?  One thing that was funny though is the beam coming directly from the laser.  There is the main, regular beam, and then there is a thin horizontal line of red light also coming straight out of the laser.  I don't know what to do about that, except perhaps put an iris right after the HeNe, to block the horizontal part?  I'm not sure that it's doing anything bad to the optic though, since the horizontal part gets clipped by other optics before the beam enters the vacuum, so there is no real irregularity on the beam incident on the QPD.

I realigned the oplev on the QPD, using last night's ITMX alignment + whatever drift it picked up over night, so it may need re-recentering after Xarm is nicely aligned.

ELOG V3.1.3-